Find the solution by substitution:
4x+3y=5
2x-y=-5

Answer by Alan3354(69443)

Find the solution by substitution:
4x+3y=5 eqn 1
2x-y=-5 eqn 2
--------------
-y = -5 - 2x (eqn 2)
y = 2x+5
Sub into eqn 1
4x + 3*(2x+5) = 5
4x + 6x+15 = 5
10x = -10
x = -1
--------
y = 2x+5
y = 3